OpenStudy (anonymous):

somebody please help me with this question. Mr. Thompson has been asked to add storage space to a garage. He estimates 6 square feet per individual will be required to meet the needs of 12 people. If the storage space is 15% of the total amount of space required, how much total space does Mr. Thompson have? Mr. Thompson had______ square feet.

OpenStudy (anonymous):

A good start is to see how much total storage space is required. So, that part is given by 6 square feet for each person and there are 12 people.

OpenStudy (anonymous):

6x12=72

OpenStudy (anonymous):

Good. Now that you have that number. The initial number that you are starting with is only 15% of that.

OpenStudy (anonymous):

72x0.15

OpenStudy (anonymous):

480

OpenStudy (anonymous):

72/0.15=480

OpenStudy (anonymous):

Yes, and you are answering your own questions beautifully, which is the goal. Except that you have to multiply out 72 by 0.15 to get 10.8. So, yes to your first of the 3 answers above.

OpenStudy (anonymous):

An alloy weighing 30 lbs. is 11% tin. The alloy was made by mixing a 15% tin alloy and a 9% tin alloy. How many pounds of each alloy were used to make the 11% alloy? ___ lbs. of the 15% alloy and ___ lbs. of the 9% alloy.

OpenStudy (anonymous):

The 15% is basically saying that he needs 72, but he only has 15% of that to start with, so by eyeballing the 72 x 0.15, we know that it will be a fairly small number. For checking.

OpenStudy (anonymous):

can you help me with this one i dont get it

OpenStudy (anonymous):

This one is a little harder, but I can definitely help you.

OpenStudy (anonymous):

A lot of problems like this are made easier by first correctly conceptualizing the problem, and then the math is actually the easier part. We start by asking how much tin is actually in the 30 lbs. In 30 lbs., we have 0.11 of it as tin, so we have 30 x 0.11 lbs. or 3.3 lbs. Believe it or not, that's half the problem done.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

30x0.15=4.5

OpenStudy (anonymous):

i got 3.3 first answer and second one i got 7.2

OpenStudy (anonymous):

Now, let's call the 15% alloy as x lbs. and the 9% alloy as y lbs. The total weight we know from the problem is 30 lbs. So, we know that x + y = 30 because that is the total weight. We also know that (x)(0.15) + (y)(0.09) = 3.3 because we know we have 3.3 lbs. of tin total. So, we have 2 equations in 2 unknowns.

OpenStudy (anonymous):

can you put the in one equation little confused

OpenStudy (anonymous):

ok what is next i got this one

OpenStudy (anonymous):

now we have to solve for x right???

OpenStudy (anonymous):

So, are you ok with the 2 equations in the 2 unknowns, or are you stuck at this point?

OpenStudy (anonymous):

i need to know this ______ lbs. of the 15% alloy and ___ lbs. of the 9% alloy.

OpenStudy (anonymous):

there is a third part to this right??? solve for x

OpenStudy (anonymous):

3.3 is the first ansnwer right.

OpenStudy (anonymous):

Yes, that's from the problem statement. You have x + y = 30 and you have (0.15)x + (0.09)y = 3.3 These are your 2 equations for going forward with the problem. What I am asking is if you understand the derivation of the equations and the steps up to this point. Plus, a second question is are you able to solve simultaneous equations. I asked 2 questions. What are your 2 answers?

OpenStudy (anonymous):

y=35

OpenStudy (anonymous):

I'm looking for 2 yes/no answers at this point. Then we can proceed.

OpenStudy (anonymous):

i dont know what you looking for sorry

OpenStudy (anonymous):

Please read my posts.

OpenStudy (anonymous):

yes i do understand

OpenStudy (anonymous):

So, then all that is left is to solve the simultaneous equations for x and y, and from your last answer, you indicate that you are able to do that.

OpenStudy (anonymous):

x=21.4

OpenStudy (anonymous):

y=35

OpenStudy (anonymous):

x=3.21 y=3.15

OpenStudy (anonymous):

Your 2 simultaneous equations are x + y = 30 and (0.15)x + (0.09)y = 3.3. Solving the first equation for x, we have x = 30 - y. We substitute that into the second equation and solve for y. Y=20. We go back to the first equation with that y and find that x=10. So, 10 lbs. of the 15% alloy and 20 lbs. of the 9% alloy. And that's your answer.

OpenStudy (anonymous):

will you tell me how you got that

OpenStudy (anonymous):

(0.15)x + (0.09)y = 3.3 -> (0.15)(30-y) + (0.09)y = 3.3. And solve for y. This is a little bit more detail flowing from my previous post.

OpenStudy (anonymous):

(0.15)(30-y) + (0.09)y = 3.3 -> 4.5 - (0.15)y + (0.09)y = 3.3. Combine y's.